OpenStudy (meilendurcer):

Find the LCD of 3/5 and 1/6. 5 6 11 30

OpenStudy (meilendurcer):

0?

OpenStudy (skullpatrol):

Not zero.

OpenStudy (skullpatrol):

Want to try again?

OpenStudy (anonymous):

Find the LCD first. 3/5 = 18/30 1/6 = 5/30 18 - 5 = 13 Answer: 13/30.

OpenStudy (anonymous):

The LCD is 30 because it is the lowest number that both 5 and 6 could go into.

OpenStudy (meilendurcer):

Thanks!

OpenStudy (anonymous):

no problem!